OpenStudy (anonymous):

A drawer contains 3 tank tops, 6 short-sleeve shirts, 4 three-quarter sleeve shirts, and 5 long-sleeve shirts. What is the probability of randomly choosing a tank top and then a three-quarter sleeve length, if you replace the tank top?

OpenStudy (anonymous):

You have a sample size of 3 + 6 + 4 + 5 objects. There are 3 tank tops. What is the probability?

OpenStudy (anonymous):

2/3 ?

OpenStudy (anonymous):

No, you have 18 total objects. You've got one shot to pick 3 out of 18 of them. What are those odds?

OpenStudy (anonymous):

\[\frac{\text{Number of tank tops}}{\text{Number of total objects}} \times \frac{\text{Number of three-quarter sleeves}}{\text{Number of total objects}}\]

OpenStudy (anonymous):

7/18 ?

OpenStudy (anonymous):

No, you aren't adding them.

OpenStudy (anonymous):

thenn 1/18

OpenStudy (anonymous):

When you multiply fractions remember \[a / b * x / y = ax/by\] \[(3/18)(4/18)\]

OpenStudy (anonymous):

12/ 18?

OpenStudy (anonymous):

What is 3 * 4?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

And 18 * 18?

OpenStudy (anonymous):

324

OpenStudy (anonymous):

And what is 12/324?

OpenStudy (anonymous):

ohhhh 1/27 ?

OpenStudy (anonymous):

Correct, or roughly 3.7%. Wouldn't put a high bet on that happening. :]